OpenStudy (anonymous):

Solve for u. -3/2-3/5u=-7/3

OpenStudy (anonymous):

Add 3/2 to both sides. -3/5u = -5/6 Multiply both sides by -5/3 u = 25/18

OpenStudy (anonymous):

-3/5u = -7/3 + 3/2 = -5/6; cross multiply: -3*6 = 5u*-5; -18 = 5u*-5; 5u = -18/-25 = 18/25

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

I am pretty sure it is 18/25, not 25/18

OpenStudy (anonymous):

Try plugging in 25/18 into the equation and see if you get -7/3 :)

OpenStudy (anonymous):

hey can you help me understand where I went wrong? I get that you are right :-D

OpenStudy (anonymous):

You put the variable in the denominator while cross multiplying rather than in the numerator.

OpenStudy (anonymous):

Oh I know what happened. I took u as being in the denominator. I depends on what the original problem was.